Randall and Amyra left school at the same time and began walking in opposite directions. Randall walked at a rate of 3.6 mph and amyra walked at a rate of 4.2 mph. How far apart were the after 10 minutes
Answer by stanbon(48526)

10 minutes is 1/6 of an hour.
Distance = rate*(time)
3.6(1/6)+4.2(1/6)= 0.6 + 0.7 = 1.3 miles
Cheers,
Stan H.